Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2014-04-29

Observers: Sandra Benitez, Joe Polshaw
Support: Kate Maguire 
TAT on duty: Andrea Pastorello, Cosimo Inserra

--------------------------------------------------------------

WEATHER CONDITIONS:
start: wind 2.3 m/s; temp. 8.2 C; humid. 14%; Clouds across the whole sky. 
middle: wind 4.8 m/s; temp. 8.4 C: humid. 13%; Thin clouds.
end: wind 5.8 m/s; temp. 6.7; humid. 14%; Clear

Complete cloud cover for the first half of the night. The clouds gradually became thinner during the second half, and had cleared by the end of the night.

--------------------------------------------------------------

INSTRUMENTS: EFOSC2

CALIBRATIONS

Afternoon:
-- EFOSC bias, arcs, dome flats

Morning:
-- EFOSC arcs

--------------------------------------------------------------

OBSERVATIONS

Target                   Type                         UT-Start         Filter/Grism           Slit (")         Seeing (")      Notes

*** cloud cover from beginning of the night ***

Clouds thin and dome opened at ~04:20 UT for image analysis

SN2014ad            Follow. (SPEC)            04:45              Gr11+16                1                1 to 2           (1)   
LSQ14asn             Follow. (SPEC)            06:07             Gr11                       1                 1                 (2)
EG274                  STD (SPEC)                 07:58             Gr11/13/16            1                0.6               (3)
MLS140425...       Class                         08:22             Gr13                        1                0.8               (4)
SNhunt242           Class                         09:12             Gr13                        1                0.8               (5)
EG131                   STD (SPEC)                09:58             Gr11/13/16            1                0.8                -      

(1) Observing through thin clouds. Seeing in the acquisition image is around 2", but the DIMM seeing is 0.6" to 1". Seeing improving so we decided to stick with the 1" slit. 1200s in gr11+16 (increased exptime from standard OB to account for clouds).
(2) Still some thin clouds. We lost time here since there was a problem focusing the telescope. 2*2700s Gr11. S/N looks quite low.
(3) Lost more time again since the image analysis needed to be re-done due to improved conditions. No clouds now.
(4) MLS140425-152525-191348. Looks faint in acquisition image. 1800s exposure. Extremely noisy spectrum...
(5) Some time lost since the wrong target was selected at first. Not sure what this is - SNID thinks it could be a Ib/c, but can't find a nice fit. Spectrum is a little bit red. Strong emission at ~7100A which looks to have a narrow and broader component.

Comments